# Paired Image Generation with Diffusion-Guided Diffusion Models

### v1

@@ Abstract
The segmentation of mass lesions in digital breast tomosynthesis (DBT) images is very significant for the early screening of breast cancer. However, the high-density breast tissue often leads to high concealment of the mass lesions, which makes manual annotation difficult and time-consuming. As a result, there is a lack of annotated data for model training. Diffusion models are commonly used for data augmentation, but the existing methods face two challenges. First, due to the high concealment of lesions, it is difficult for the model to learn the features of the lesion area. This leads to the low generation quality of the lesion areas, thus limiting the quality of the generated images. Second, existing methods can only generate images and cannot generate corresponding annotations, which restricts the usability of the generated images in supervised training. In this work, we propose a paired image generation method. The method does not require external conditions and can achieve the generation of paired images by training an extra diffusion guider for the conditional diffusion model. During the experimental phase, we generated paired DBT slices and mass lesion masks. Then, we incorporated them into the supervised training process of the mass lesion segmentation task. The experimental results show that our method can improve the generation quality without external conditions. Moreover, it contributes to alleviating the shortage of annotated data, thus enhancing the performance of downstream tasks.
Keywords:
Paired image generationDiffusion-guided diffusion models DBT mass segmentation
@@ 2.2 Paired Image Generation
The probability distribution $p_{\theta}$ of the joint denoising process satisfies Eq. (11), which can be implemented with two diffusion models $x$ and $y$ to fit $p_{x}$ and $p_{y}$ respectively as proved in Eq. (2.2.2). In Eq. (11)-(2.2.2), $P(\cdot)$ is the abbreviation of probability distribution, $P(x_{T})=\mathcal{N}(x_{T};0,I)$ and $P(y_{T})=\mathcal{N}(y_{T};0,I)$. Specifically, we can train an additional diffusion guider as model $x$. It takes $y_{t}$ as the guiding signal to generate $x_{t-1}$ for guiding model $y$. Moreover, according to Eq. (6)-(7), in the generation process of $x_{t-1}$, $\hat{x}_{0}$ will be predicted first. Therefore, we can use the conditional diffusion model $p_{y}(y_{t-1}|x_{0},y_{t})$ as model $y$ to avoid the influence of the known noise $\epsilon_{x}$. During the generation process, this diffusion-guided diffusion model architecture can generate guiding signals for each other, thereby improving the generation quality. The training and sampling processes are shown in Alg 1-2, where we set $\sigma_{t}=0$.
@@
4:     $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{1},y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{2}$triangleright Eq. (9)-(10)
5:     $\epsilon_{x}=model_{x}(x_{t},y_{t},t)$triangleright $p_{x}(x_{t-1}|x_{t},y_{t})$
@@
4:     $\epsilon_{x}=model_{x}(x_{t},y_{t},t)$triangleright $p_{x}(x_{t-1}|x_{t},y_{t})$
5:     $\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{x}}{\sqrt{\bar{\alpha}_{t}}},x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\epsilon_{x}$triangleright Eq. (6)-(7)
6:     $\epsilon_{y}=model_{y}(\hat{x}_{0},y_{t},t)$triangleright $p_{y}(y_{t-1}|\hat{x}_{0},y_{t})=p_{y}(y_{t-1}|\frac{x_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{x}}{\sqrt{\bar{\alpha}_{t-1}}},y_{t})$
7:     $\hat{y}_{0}=\frac{y_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{y}}{\sqrt{\bar{\alpha}_{t}}},y_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{y}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\epsilon_{y}$triangleright Eq. (6)-(7)
@@ 3.1 Datasets and Experimental Setup
This private DBT mass segmentation dataset is sourced from **** Hospital and **** Hospital, and its ethics application number is ****. The dataset contains data of 367 patients and the data of each patient includes images from the CC and MLO perspectives and the corresponding mass lesion masks, which are manually annotated by two experienced radiologists. There are a total of 8,723 slices containing mass lesions. In the preprocessing stage, we cropped the blank parts and resized the images to 512×512.
@@
We linearly scaled the images to $[-1,1]$ and set the noise schedule $\beta_{t}$ to constants increasing linearly from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$ [5]. The maximum time step $T$ was set to 1024. We used U-Net [14] as the network architecture of the two diffusion models. We set the number of the input channels to 2 and introduced the guiding signal through concatenation in the channel dimension [9, 7]. During the sampling process, we adopted the uniform step sampling method of DDIM [15], with the number of sampling steps set to 256.
@@ 3.2 Results
Since the proposed PIG method is a generation approach that does not require external conditions, we conducted comparative experiments with the unconditional diffusion models, including DDPM [5] and DDIM [15]. Specifically, we set the number of sampling steps to 1024 for DDPM and 256 for DDIM, and employed each method to generate 2048 DBT slices respectively. We adopted the Frechet Inception Distance (FID) [3] metric to evaluate the quality of the generated images.
@@
The experimental results are presented in Table 1. As can be seen from the results, by introducing the diffusion guidance, the network can focus more on generating the lesion parts. Therefore, the quality of the generated images is enhanced without external conditions.
@@
Dataset Dice(%$\uparrow$ IoU(%$\uparrow$ Precision(%$\uparrow$ Recall(%$\uparrow$
@@
The average results of the five-fold cross-validation experiment are presented in Table 2. Here, "+PIG_2048" denotes the addition of 2048 annotated images generated by the PIG method to the original DBTMassSeg dataset and the notations of the other datasets follow the same principle. As can be observed from the results, since the PIG method can generate more diverse lesion masks, when generating the same number of DBT slices, the images generated by the PIG method can help the network learn the features of lesion areas better, thereby enhancing the network’s generalization and performance.
@@ 4 Conclusion
We propose an innovative unconditional paired image generation method. First, through mathematical derivations, we prove that the generation process of paired images can be achieved by two diffusion processes guided mutually. Therefore, by training an additional diffusion guider for the conditional diffusion model, we can introduce the guiding signal into the diffusion process without external conditions. The experimental results show that our generation method can improve the generation quality and effectively alleviate the shortage of annotated data.

### v2

@@ Abstract
The segmentation of mass lesions in digital breast tomosynthesis (DBT) images is very significant for the early screening of breast cancer. However, the high-density breast tissue often leads to high concealment of the mass lesions, which makes manual annotation difficult and time-consuming. As a result, there is a lack of annotated data for model training. Diffusion models are commonly used for data augmentation, but the existing methods face two challenges. First, due to the high concealment of lesions, it is difficult for the model to learn the features of the lesion area. This leads to the low generation quality of the lesion areas, thus limiting the quality of the generated images. Second, existing methods can only generate images and cannot generate corresponding annotations, which restricts the usability of the generated images in supervised training. In this work, we propose a paired image generation method. The method does not require external conditions and can achieve the generation of paired images by training an extra diffusion guider for the conditional diffusion model. During the experimental phase, we generated paired DBT slices and mass lesion masks. Then, we incorporated them into the supervised training process of the mass lesion segmentation task. The experimental results show that our method can improve the generation quality without external conditions. Moreover, it contributes to alleviating the shortage of annotated data, thus enhancing the performance of downstream tasks. The source code is available at https://github.com/zhanghx1320/PIG.
@@ 2.2 Paired Image Generation
The probability distribution $p_{\theta}$ of the joint denoising process satisfies Eq. (11), which can be implemented with two diffusion models $x$ and $y$ to fit $p_{x}$ and $p_{y}$ respectively as proved in Eq. (2.2.2). In Eq. (11)-(2.2.2), $P(\cdot)$ is the abbreviation of probability distribution, $P(x_{T})=\mathcal{N}(x_{T};0,I)$ and $P(y_{T})=\mathcal{N}(y_{T};0,I)$. Specifically, we can train an additional diffusion guider as model $x$. It takes $y_{t}$ as the guiding signal to generate $x_{t-1}$ for guiding model $y$. Moreover, according to Eq. (6)-(7), in the generation process of $x_{t-1}$, $\hat{x}_{0}$ will be predicted first. Therefore, the conditional diffusion model $p_{y}(y_{t-1}|x_{0},y_{t})$ (such as Maisi [2], a generative foundation model pretrained on large-scale unannotated datasets) can be used as model $y$ to avoid the influence of the known noise $\epsilon_{x}$. During the generation process, this diffusion-guided diffusion model architecture can generate guiding signals for each other, thereby improving the image quality. The training and sampling processes are shown in Alg 1-2, where we set $\sigma_{t}=0$.
@@
4:  $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{1},y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{2}$ $\triangleright$ Eq. (9)-(10)
5:  $\epsilon_{x}=model_{x}(x_{t},y_{t},t)$ $\triangleright$ $p_{x}(x_{t-1}|x_{t},y_{t})$
@@
4:  $\epsilon_{x}=model_{x}(x_{t},y_{t},t)$ $\triangleright$ $p_{x}(x_{t-1}|x_{t},y_{t})$
5:  $\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{x}}{\sqrt{\bar{\alpha}_{t}}},x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\epsilon_{x}$ $\triangleright$ Eq. (6)-(7)
6:  $\epsilon_{y}=model_{y}(\hat{x}_{0},y_{t},t)$ $\triangleright$ $p_{y}(y_{t-1}|\hat{x}_{0},y_{t})=p_{y}(y_{t-1}|\frac{x_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{x}}{\sqrt{\bar{\alpha}_{t-1}}},y_{t})$
7:  $\hat{y}_{0}=\frac{y_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{y}}{\sqrt{\bar{\alpha}_{t}}},y_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{y}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\epsilon_{y}$ $\triangleright$ Eq. (6)-(7)
@@ 3.1 Datasets and Experimental Setup
This private DBT mass segmentation dataset is sourced from Suzhou Municipal Hospital and Guizhou Provincial People’s Hospital [^1] The data used in this study were approved by the Ethics Committee of Suzhou Municipal Hospital (No. 2024320) and Guizhou Provincial People’s Hospital (No. 2024328). Written informed consent was waived as the data were fully anonymized.. The dataset contains data of 367 patients and the data of each patient includes images from the CC and MLO perspectives and the corresponding mass lesion masks, which are manually annotated by two experienced radiologists. There are a total of 8,723 slices containing mass lesions. In the preprocessing stage, we cropped the blank parts and resized the images to 512×512.
@@
We linearly scaled the images to [1,1][-1,1] and set the noise schedule $\beta_{t}$ to constants increasing linearly from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$ [5]. The maximum time step $T$ was set to 1024. We used U-Net [14] as the network architecture of the two diffusion models. We set the number of the input channels to 2 and introduced the guiding signal through concatenation in the channel dimension [9, 7]. During the sampling process, we adopted the uniform step sampling method of DDIM [15], with the number of sampling steps set to 256.
@@ 3.2 Results
Since the proposed PIG method is a generation approach that does not require external conditions, we conducted comparative experiments with the unconditional diffusion models, including DDPM [5] and DDIM [15]. Specifically, we set the number of sampling steps to 1024 for DDPM and 256 for DDIM, and employed each method to generate 2048 DBT slices respectively. We adopted the Frechet Inception Distance (FID) [3] metric to evaluate the generation quality.
@@
The experimental results presented in Table 1 show that during the image generation process, PIG generates guiding signals simultaneously, which ensures that the process requires no additional input conditions and enhances the image generation quality.
@@
[…]
@@
The average results of the five-fold cross-validation experiment are presented in Table 2. Here, "+PIG_2048" denotes the addition of 2048 annotated images generated by the PIG method to the original DBTMassSeg dataset and the notations of the other datasets follow the same principle. As shown in the results, since PIG can generate more diverse lesion masks, when generating the same number of DBT slices, the images generated by PIG can help the network learn the features of lesion areas better, thereby enhancing the network’s generalization and performance. Additionally, in the experiment, all the generation methods exhibit similar trends. Specifically, the performance peaks after generating about 2048 images, and further increasing the proportion of generated images leads to performance fluctuations. For PIG, the performance dips to its lowest point after generating about 3072 images, which outperforms the highest values achieved by other methods. This indicates the high quality and stability of PIG.
@@ 4 Conclusion
We propose an innovative unconditional paired image generation method. First, through mathematical derivations, we prove that the generation process of paired images can be achieved by two diffusion processes guided mutually. Therefore, by training an additional diffusion guider for the conditional diffusion model, we can introduce the guiding signal into the diffusion process without external conditions. The experimental results show that our generation method can improve the generation quality and effectively alleviate the shortage of annotated data. {credits}
4.0.1 Acknowledgements
This study was partly supported by National Natural Science Foundation of China (No. 62371449 and No. 82302286), Jiangsu Provincial Key Research and Development Program Social Development Project (No. BE2022720), Guizhou Provincial Basic Research Program (Natural Science) (zk[2025]-513), and Guizhou Provincial Health Commission Science Technology Fund Project (gzwkj2024-474).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.